# SynHLMA: Synthesizing Hand Language Manipulation for Articulated Object with Discrete Human Object Interaction Representation

### v1

@@ Abstract
Generating hand grasps with language instructions is a widely studied topic that benefits from embodied AI and VR/AR applications. While transferring into Hand Articulatied Object Interaction (HAOI), the hand grasps synthesis requires not only object functionality but also long-term manipulation sequence along the object deformation. This paper proposes a novel HAOI sequence generation framework SynHLMA, to Synthesize Hand Language Manipulation for Articulated objects. Given a complete point cloud of an articulated object, we utilize a discrete HAOI representation to model each hand object interaction frame. Along with the natural language embeddings, the representations are trained by an HAOI Manipulation Language Model to align the grasping process with its language description in a shared representation space. A joint-aware loss is employed to ensure hand grasps follow the dynamic variations of articulated object joints. In this way, our SynHLMA achieves three typical hand manipulation tasks for articulated objects of HAOI generation, HAOI prediction and HAOI interpolation. We evaluate SynHLMA on our built HAOI-lang dataset and experimental results demonstrate the superior hand grasp sequence generation performance comparing with state-of-the-art. We also show a robotics grasp application that enables dexterous grasps execution from imitation learning using the manipulation sequence provided by our SynHLMA. Our codes and datasets will be made publicly available.
@@ 1 Introduction
Human-object interaction is widely studied, yet inferring human manipulation intent remains challenging—especially for enabling robots to learn dexterous manipulation through object functionality understanding and action execution. Current works primarily synthesize rigid-object hand-object interactions (HOI) with physically realistic grasp poses. However, articulated objects demand modeling complete deformation processes beyond simple grasps. For instance, scissor usage requires synthesizing both grasp poses and subsequent manipulation sequences (e.g., opening/closing motions). Thus, learning to generate Human Articulated Object Interactions (HAOI) under human intent is critical for robotic dexterity.
@@
Existing grasp synthesis approaches face significant limitations. Methods using robotic hands (Jin et al., 2024; Xie et al., 2023) lack human-hand realism, while skeleton-driven techniques (Yang et al., 2022a; Karunratanakul et al., 2021) ignore hand-object contact physics. Contact prediction models (Grady et al., 2021; Yang et al., 2022b) struggle to integrate language descriptions with articulated object dynamics. More critically, diffusion models exhibit poor long-sequence generation due to insufficient priors, and most methods neglect model generalization beyond text-to-HOI tasks.
To address the aforementioned issues, in this paper, we propose SynHLMA, a novel HAOI sequence generation framework to Synthesize Hand Language Manipulation for Articulated objects. Given an articulated object’s point cloud and language query, SynHLMA leverages VQ-VAE to discretize per-frame HAOI representations. These representations are aligned with language embeddings via LoRA-trained manipulation language models in a shared semantic space. An autoregressive strategy predicting incremental differences addresses long-sequence generation, enabling HAOI generation, prediction, and interpolation (Fig. 1).
Inspired by language’s inherent discreteness and grasp taxonomies (Feix et al., 2015), we design discrete HAOI representations with separate codebooks encoding hand position, pose, adjustments, and object configuration. During VQ-VAE reconstruction, semantic HAOI tokens are assigned per frame, supervised by our novel articulation-aware loss enforcing: hand-object penetration avoidance, pose consistency, and joint-configuration accuracy. This allows the manipulation language model to capture semantic nuances for diverse prompted tasks.
Accompanying the HAOI synthesis task, we introduce a task-specific dataset, HAOI-Lang. Leveraging a physics-based interaction engine, we sample 200 contact points and directions per articulated object to generate extensive HAOI interaction sequences. Following an initial manual filtering, GPT-4 is employed to produce diverse natural language annotations for each interaction, encompassing grasping intent, direction, and location. We evaluate SynHLMA on three tasks: HAOI generation, prediction, and interpolation. Experimental results demonstrate that SynHLMA outperforms existing hand-manipulation generation baselines. Furthermore, we show that the HAOI sequences generated by SynHLMA can effectively guide imitation learning for dexterous robotic hands. In summary, our work makes the following key contributions:
1) HAOI Language Dataset: We construct a novel dataset for articulated object grasping, which includes detailed natural language descriptions of grasp intents and actions. 2) Discrete Manipulation Learning: We discretize manipulation trajectories using hierarchical grasp tokens, improving generation quality and control. 3) Articulation-Aware Loss:By introducing a set of complementary constraints on the HAOI generation process, we mitigate distributional discrepancies among different network components and improve the model’s responsiveness to articulated-object variations. 4) Manipulation Language Model: We present the first language model for articulated object manipulation, bridging natural language and high-level actions via grasp tokenization to support HAOI generation, prediction, and interpolation.
@@ 2.1 Hand Grasp Synthesis
Probabilistic grasp generation can be broadly categorized into two approaches: regression-based and diffusion-based methods. In regression-based models (Li et al., 2024; Huang et al., 2025; Prokudin et al., 2019; Yang et al., 2022b; 2021), a point cloud encoder is first used to extract features representing the HOI. The model then learns the distribution of these features and samples from it, followed by a regression module that predicts the hand parameters to complete the grasp generation. In contrast, diffusion-based methods (Ye et al., 2023; Zhang et al., 2024b) gradually add noise to existing HOI data and learn to reverse this process through denoising. By modeling the denoising trajectory, the model captures the underlying distribution of HOI configurations in the dataset. Grasp generation is then achieved by iteratively denoising a noise-initialized sample.
@@
2.2 Multimodel Large Language Models
In recent years, the remarkable success of large language models (LLMs) such as GPT (Achiam et al., 2023), Qwen (Bai et al., 2023), and Gemini (Team et al., 2023) in various downstream natural language tasks—including text translation (Devlin et al., 2019), semantic understanding (Du et al., 2022), and text generation (Zhang et al., 2020) has inspired advancements in other domains. Many researchers have attempted to integrate LLMs with multimodal data, including audio (Borsos et al., 2023), images (Radford et al., 2021), videos (Li et al., 2023), and 3D object models (Jiang et al., 2023), to develop powerful multimodal models.
@@ 3.1 Overview
We introduce SynHLMA, a framework for generating HAOI sequences. It discretizes full grasping trajectories into token sequences via a multi-stage VQ-VAE with articulation-aware constraints. A manipulation language model learns the distribution of tokenized actions conditioned on language descriptions, enabling complete HAOI sequence generation. We construct a large-scale physics-simulated dataset, with each interaction annotated using GPT.
3.2 Synthesis HAOI Data Generation
@@
To ensure consistent data generation, we assign each object category a specific manipulation-angle range and uniformly sample grasping frames from it. For any grasp point p{p} relative to the object’s center of mass $g_{\mathcal{O}}$, the grasp direction is ${D}={p}-{g}_{\mathcal{O}}$. Each frame provides a homogeneous transformation $\mathcal{T}\in\mathbb{R}^{4\times 4}$; we normalize coordinates by applying $\mathcal{T}^{-1}\mathbf{x}$, mapping all grasping data into a unified canonical space.
@@
To capture human intent, we render GraspXL grasping sequences in Open3D and prompt GPT-4 to produce multi-perspective textual descriptions covering contact points, wrist orientation, and intended action. This ensures concise, consistent annotations (Figure 4). Our dataset includes 256 instances with over 500,000 static grasps and 50,000 manipulation sequences.
3.3 Discrete Articulated Manipulation Representation
Similar to previous works, we parameterize the grasp configuration as: $\bm{\beta}=({R},{P},{T}),$ where ${R}\in\mathbb{R}^{3}$ denotes the global rotation of the hand, $P\in\mathbb{R}^{90}$} represents the hand pose, and $T\in\mathbb{R}^{3}$ denotes the global translation of the hand. We then employ a parametric hand model—specifically, the MANO model—to render the hand mesh $\mathcal{H}\in\mathbb{R}^{778\times 3}$} based on these parameters: $\mathbf{\mathcal{H}}=\mathcal{M}(\bm{\beta})=\mathcal{M}({R},{P},{T}),$ where $\mathcal{M}(\cdot)$ denotes the MANO hand model function that maps the grasp parameters to a 3D hand mesh. In this work, to better fit the motion characteristics of the articulated object $\mathcal{O}$ and the hand during the grasping process, we design two separate VQ-VAE discretization frameworks for them, as the image 2.
To ensure consistency between articulated object motions and hand movements, as well as the physical plausibility of grasps, we introduce a joint token $j\in\mathbb{N}$ and train it using a single-stage VQ-VAE. Given an object $\mathcal{O}$ and a grasp contact point $m_{\text{contact}}$, we first extract joint-relevant features using a PointNet encoder. These features are then processed by a joint encoder $\mathcal{E}_{j}$ and concatenated with the canonical object feature $C_{\mathcal{O}}$. The combined representation is passed to a joint prediction decoder $\mathcal{D}_{\text{j}}$, which estimates the joint parameters $J\in\mathbb{R}^{3}$ (e.g., rotation angle or translation). We further introduce a joint-aware loss, which guides the subsequent manipulation generation and comprises three components:
HAOI Penetration Loss. This term penalizes interpenetration between the predicted hand mesh $\hat{\mathcal{H}}$ and the object surface. Let $V$ denote the vertices of $\hat{\mathcal{H}}$, and $V_{\text{in}}\subseteq V$ be the subset that penetrates the object mesh. For each $v_{i}\in V_{\text{in}}$, we compute its squared Euclidean distance to the closest point $v_{i}^{\mathcal{O}}$ on the object mesh. Averaging over all penetrating vertices mitigates the effect of outliers:
$\mathcal{L}_{P}=\frac{1}{|V_{\text{in}}|}\sum_{v_{i}\in V_{\text{in}}}\left\|v_{i}-v_{i}^{\mathcal{O}}\right\|_{2}^{2}$ (1)
Pose Consistency Loss. To ensure that the poses generated by the hand and object VQ-VAE architectures are aligned in both rotation and translation, we define a unified pose consistency loss as:
$\mathcal{L}_{C}=\begin{cases}\arccos\left(\frac{\mathrm{Tr}(\mathcal{R}_{J}\cdot\mathcal{R}_{R}^{\top})-1}{2}\right)\cdot\frac{180^{\circ}}{\pi},&\text{if joint is rotational,}\\[4.2679pt] \|{J}-{T}\|_{2},&\text{if joint is translational.}\end{cases}$ (2)
When predicting the rotation axis of the object, $J$ represents the axis–angle, both $J$ and $R$ are converted into rotation matrices $\mathcal{R}_{J}$ and $\mathcal{R}_{R}\in\mathbb{R}^{3\times 3}$ via the Rodrigues formula. For translational joints, $J$ represents the object displacement vector. This formulation ensures that each joint contributes an appropriate error term depending on its type.
Joint Aware Loss. We employ this loss to ensure that the VQ-VAE decoder accurately captures the angular state of the object at each frame. Here, $J_{\text{gt}}$ denotes the ground-truth values.
$\mathcal{L}_{J}=\left\|\mathcal{D}_{j}\left(\left[\mathcal{E}_{j}\left(\text{PointNet}(\mathcal{O},m_{\text{contact}})\right),\,C_{\mathcal{O}}\right]\right)-J_{gt}\right\|_{2}^{2}$ (3)
Combining these three terms, the total articulation-aware loss is defined as:
$\mathcal{L}_{\text{artic}}=\lambda_{P}\cdot\mathcal{L}_{P}+\lambda_{C}\cdot\mathcal{L}_{C}+\lambda_{J}\cdot\mathcal{L}_{J}$ (4)
For the hand, to achieve more detailed fitting of its features, we divide the hand generation process into three stages. These stages learn the global pose $\bm{\langle g\rangle}$, local articulation $\mathbf{\langle l\rangle}$, and refinement $\mathbf{\langle r\rangle}$ tokens, where ${g},{l},{r}\in\mathbb{N}$. We utilize a multi-stage VQ-VAE architecture, which consists of three pairs of encoders $\mathcal{D}_{i}$, decoders $\mathcal{E}_{i}$, and codebooks $\mathcal{B}_{i}$, where $i\in\{{g},{l},{r}\}$. The hand’s point cloud features and the object’s discretized features are input into this architecture. After discretization, the features are reconstructed via the corresponding decoders to generate the final grasping configuration.
[…]
In our framework, for the codebooks $\mathcal{B}_{j}$ and $\mathcal{B}_{i},\;i\in\{{g},{l},{r}\}$, each feature vector output $Q$ by the encoder is compared to the discrete vectors $q$ in the codebook by computing the squared Euclidean distance. Each feature vector is then replaced with its nearest discrete codebook vector. Assume that each codebook contains $k$ vectors, each of dimension $d$. This quantization process can be formulated as:
${q_{i}}$ $=\mathcal{E}_{i}({Q}),\quad i\in\{g,l,r,j\}$ (5)
index $=\arg\min_{k}\|{q}-{b}_{t}\|^{2},\quad t\in\{0,1,\ldots,k-1\}$
$\quad{q_{i}}$ $=\mathcal{B}_{i}[\text{index}].$
VQ-VAE training involves two primary losses: reconstruction loss, which drives the decoder to match targets, and commitment loss, which keeps encoder outputs close to the codebook while promoting latent-space diversity. The reconstruction loss for the single-stage VQ-VAE applied to articulated objects is already accounted for in Equation 3, so only the commitment loss is computed. The overall loss is obtained by summing all individual terms, formally defined below:
$\mathcal{L}_{\text{r}}$ $=\lambda_{1}\cdot\|\mathcal{M}(R,0,T)-\mathcal{M}(\hat{R},0,\hat{T})\|^{2}$ (6)
$+\lambda_{2}\cdot\|\mathcal{M}(R,P,T)-\mathcal{M}(\hat{R},\hat{P},\hat{T})\|^{2}$
$+\lambda_{3}\cdot\|\mathcal{M}(R,P,T)-\mathcal{M}(\Delta\hat{R}\times\hat{R},\Delta\hat{P}\times\hat{P},\Delta\hat{T}+\hat{T})\|^{2}$
Meanwhile, the commit loss for tokens at each layer can be expressed by the following:
$\mathcal{L}_{\text{c}}=\sum_{i\in\{g,l,r,j\}}\|\text{sg}[{q_{i}}]-\mathcal{N}_{i}(q_{i})\|^{2}+\beta\|{q}-\text{sg}[\mathcal{N}_{i}(q_{i})]\|^{2}$ (7)
where $\text{sg}[\cdot]$ denotes the stop-gradient operator , $\beta$ is a hyperparameter balancing the commitment cost and $\mathcal{N}_{i}(\cdot)$ denotes the discretized value in the codebook of token $i$ that is closest to the feature $q_{i}$. The goal of VQ-VAE is to minimize $\mathcal{L}_{\text{artic}}$, $\mathcal{L}_{\text{r}}$ and $\mathcal{L}_{\text{c}}$ as much as possible.
3.4 HAOI Manipulation Language Model
To train a HAOI manipulation language model for grasp perception, we first extract the target object’s point cloud using PointNet. As the text primarily describes how to operate the object (e.g., shutting down a computer) including directions and positions, we do not directly provide the contact-point heatmap at this stage. Instead, we obtain the discretized joint token $\langle j\rangle$ using the pre-trained encoder $\mathcal{E}_{j}$. Each frame’s $\langle j\rangle$ token is inserted between the corresponding $\langle SG\rangle$ (Start of Grasp) and $\langle EG\rangle$ (End of Grasp) markers. The resulting sequence is concatenated with the output of the text tokenizer and fed into the model. Depending on the downstream task, the model can also accept the full HAOI representation $\langle g,l,r,j\rangle$, as illustrated in Figure 4.
We represent the manipulation process as a sequence of static frames, each discretized into hierarchical tokens for global motion $\langle g\rangle$, local motion $\langle l\rangle$, refinement $\langle r\rangle$, and object joints $\langle j\rangle$. Thus, each frame is represented as the composite token sequence $\langle g\rangle\langle l\rangle\langle r\rangle\langle j\rangle$, and the full grasping sequence can be expressed as:
$\langle HO\rangle\langle g_{1}\rangle\langle l_{1}\rangle\langle r_{1}\rangle\langle j_{1}\rangle\cdots\langle g_{t}\rangle\langle l_{t}\rangle\langle r_{t}\rangle\langle j_{t}\rangle\langle HO\rangle,$
where $\langle HO\rangle$ denotes the start and end of each frame, and $t$ is the total number of steps. For HAOI manipulation generation, $\langle HO\rangle$ tokens are removed, and each tuple $\langle g,l,r,j\rangle$ is sequentially mapped through the pre-trained multi-layer VQ-VAE codebook to obtain latent vectors. These vectors are then decoded by the multi-layer decoder to produce the grasp configuration for each frame.
$P(x_{i}|x_{<i})=P(x_{i}|x_{1},x_{2},\ldots,x_{i-1}),i=1,2,...,t.$ (8)
with the objective of minimizing the negative log-likelihood loss:
$\mathcal{L}_{\text{NLL}}=-\sum_{i=1}^{t}\log P(x_{i}|x_{<i})$ (9)

### v2

@@ Abstract
Generating language-guided hand grasps has been widely studied in embodied AI. However, extending grasp synthesis to Hand Articulated Object Interaction (HAOI) requires modeling not only object functionality but also temporally coherent manipulation along object articulation. This paper proposes a novel HAOI sequence generation framework SynHLMA, to Synthesize Hand Language Manipulation for Articulated objects. Given an articulated object point cloud and a textual instruction, SynHLMA learns a hierarchical discrete representation of hand–object interaction frames and aligns them with language embeddings through a manipulation language model in a shared semantic space. An articulation-aware objective further enforces geometric validity and joint-consistent temporal dynamics, enabling physically grounded long-horizon generation. SynHLMA supports HAOI generation, prediction, and interpolation within a single autoregressive formulation. We evaluate our method on the newly constructed HAOI-Lang dataset and demonstrate superior performance over state-of-the-art baselines. Moreover, the generated manipulation sequences effectively facilitate imitation learning for dexterous robotic grasp execution. Code and dataset will be publicly released.
Machine Learning, ICML
@@ 1 Introduction
Understanding human manipulation intent in human–object interaction (HOI) remains a fundamental challenge for robotic dexterity. While recent advances have enabled physically plausible grasp synthesis for rigid objects, articulated objects require modeling not only stable grasp poses but also temporally coherent deformation processes. For example, using scissors entails both grasp acquisition and coordinated articulation control (e.g., opening and closing). Such interactions demand reasoning over evolving object configurations and state-dependent affordances. We refer to this problem as Human Articulated Object Interaction (HAOI) generation under language-guided intent.
@@
Despite substantial progress in grasp synthesis, existing approaches exhibit notable limitations. Robotic-hand-based methods (Jin et al., 2024; Xie et al., 2023) often lack human-hand realism, while skeleton-driven techniques (Yang et al., 2022a; Karunratanakul et al., 2021) neglect physically grounded contact modeling. Contact-centric approaches (Grady et al., 2021; Yang et al., 2022b) struggle to integrate language semantics with articulated object dynamics. Moreover, diffusion-based models suffer from degraded long-horizon generation due to weak structural priors, and most prior work focuses narrowly on text-to-grasp tasks without modeling full manipulation sequences.
Articulated objects introduce intrinsic structural variability into HOI. Unlike rigid manipulation—where contact topology and kinematic constraints remain largely stable—articulated interactions exhibit configuration-dependent contacts, joint-coupled motion dependencies, and temporally evolving affordances. These properties frequently induce inconsistencies between hand motion and object deformation, significantly increasing the difficulty of coherent sequence generation.
To address these challenges, we propose SynHLMA, a unified framework for Synthesizing Hand–Language Manipulation of Articulated objects. Our key insight is that articulated manipulation exhibits discrete structural regularities analogous to linguistic tokens and grasp taxonomies (Feix et al., 2015). We therefore introduce a hierarchical discrete representation for HAOI. Specifically, two modular VQ-VAE models encode object articulation states, global hand configuration, local pose parameters, and refinement residuals into structured manipulation tokens. This design enables semantic alignment between articulation states and grasp configurations, promoting coordinated hand–object dynamics across varying object states.
Building upon the discrete representation, we develop a manipulation language model that aligns tokenized HAOI sequences with language embeddings via LoRA adaptation in a shared semantic space. An autoregressive formulation predicts incremental state differences, improving long-horizon stability and enabling three tasks: HAOI generation, prediction, and interpolation (Fig. 1).
A central challenge in articulated manipulation modeling lies in preserving articulation consistency across geometry, kinematics, and temporal evolution. Existing approaches typically optimize either continuous pose reconstruction or discrete token likelihood, without explicitly enforcing articulation constraints. Consequently, generated sequences may exhibit interpenetration, inconsistent joint states, or temporally incoherent motion.
To overcome this limitation, we introduce an articulation-aware objective that integrates multi-level structural constraints into training. Specifically, our objective enforces: (1) geometric validity to prevent hand–object penetration and ensure accurate joint decoding; (2) hierarchical reconstruction consistency between discrete tokens and continuous kinematics; (3) temporal articulation coherence across sequential states. By embedding articulation awareness directly into representation learning and sequence modeling, SynHLMA generates physically grounded and semantically consistent manipulation trajectories.
To support this task, we construct HAOI-Lang, a large-scale dataset of articulated manipulation sequences with language annotations. Using a physics-based interaction engine, we sample diverse contact points and articulation trajectories for each object. GPT-4 is employed to generate rich textual descriptions covering intent, direction, and spatial relations, followed by human refinement to ensure semantic fidelity. Experiments demonstrate that SynHLMA consistently outperforms prior baselines on HAOI generation, prediction, and interpolation, and that the synthesized sequences effectively facilitate imitation learning for dexterous robotic hands.
In summary, our work makes the following key contributions: 1) Discrete Manipulation Representation. We propose a hierarchical tokenization scheme for articulated manipulation, enabling structured and controllable sequence generation. 2) Manipulation Language Model. We present a language-aligned generative model for articulated object manipulation supporting generation, prediction, and interpolation. 3) Articulation-Aware Objective. We introduce a unified training objective that enforces geometric validity, joint-state alignment, and temporal coherence. 4) HAOI-Lang Dataset. We construct a new language-annotated dataset for articulated hand–object interaction.
@@ 2.1 Hand Grasp Synthesis
Probabilistic grasp generation can be broadly categorized into two approaches: regression-based and diffusion-based methods. In regression-based models (Li et al., 2024; Huang et al., 2025; Prokudin et al., 2019; Yang et al., 2022b, 2021), a point cloud encoder is first used to extract features representing the HOI. The model then learns the distribution of these features and samples from it, followed by a regression module that predicts the hand parameters to complete the grasp generation. In contrast, diffusion-based methods (Ye et al., 2023; Zhang et al., 2024b) gradually add noise to existing HOI data and learn to reverse this process through denoising. By modeling the denoising trajectory, the model captures the underlying distribution of HOI configurations in the dataset. Grasp generation is then achieved by iteratively denoising a noise-initialized sample.
@@
2.2 Multimodal Large Language Models
In recent years, the remarkable success of large language models (LLMs) such as GPT (Achiam et al., 2023), Qwen (Bai et al., 2023), Gemini (Team et al., 2023) and Gemma (Team et al., 2024) in various downstream natural language tasks—including text translation (Devlin et al., 2019), semantic understanding (Du et al., 2022), and text generation (Zhang et al., 2020) has inspired advancements in other domains. Many researchers have attempted to integrate LLMs with multimodal data, including audio (Borsos et al., 2023), images (Radford et al., 2021), videos (Li et al., 2023), and 3D object models (Jiang et al., 2023), to develop powerful multimodal models.
@@ 3.1 Overview
We propose SynHLMA, a unified framework for generating HAOI sequences. SynHLMA employs two modular VQ-VAE models to discretize articulated object joint trajectories and hand grasps into structured tokens, upon which a manipulation language model learns action distributions conditioned on language instructions. To ensure physically plausible and structurally consistent generation, we introduce an articulation-aware objective that enforces geometric validity, joint-state alignment, and temporal coherence. Furthermore, we construct a large-scale physics-based simulation dataset with GPT-assisted annotations and manual refinement.
3.2 Discrete Articulated Manipulation Representation
We represent a hand grasp configuration as $\boldsymbol{\beta}=(R,P,T),$ where $R\in\mathbb{R}^{3}$ denotes the axis–angle representation of the global hand rotation, $P\in\mathbb{R}^{90}$ represents the articulated hand pose parameters, and $T\in\mathbb{R}^{3}$ denotes the global hand translation. Given $\boldsymbol{\beta}$, the MANO parametric hand model $\mathcal{M}(\cdot)$ produces the corresponding hand mesh $\mathcal{H}\in\mathbb{R}^{778\times 3}$$\mathbf{\mathcal{H}}=\mathcal{M}(\boldsymbol{\beta})=\mathcal{M}({R},{P},{T})$.
Unlike rigid-object manipulation, articulated objects introduce configuration-dependent geometric and kinematic variations, which require explicit modeling of joint states and their interaction with hand motion. To achieve semantically structured and articulation-consistent representations, we discretize both object articulation and hand manipulation into tokenized latent spaces using two modular VQ-VAE models. (Fig. 2).
We first discretize the articulated object state. Let $J\in\mathbb{R}^{3}$ denote the object joint parameters (e.g., rotational or translational state). We employ a single-layer VQ-VAE with encoder $\mathcal{E}_{j}$, decoder $\mathcal{D}_{j}$, and codebook $\mathcal{B}_{j}$ to quantize $J$ into a discrete token $\langle j\rangle$, where $j\in\mathbb{N}$. This mapping converts continuous joint variations into semantically meaningful indices, from which the original joint parameters can be reconstructed. By explicitly tokenizing articulation states, we establish a compact and structured representation of object configuration that can condition downstream manipulation generation.
To model the structured nature of human grasping, we decompose manipulation into three hierarchical components: $\langle g,l,r\rangle,$ corresponding to global hand configuration, local pose articulation, and refinement residuals, respectively. Each component is quantized using a dedicated VQ-VAE branch with codebook $\mathcal{B}_{i}$, encoder $\mathcal{E}_{i}$, and decoder $\mathcal{D}_{i}$, where $i\in\{g,l,r\}$.
This hierarchical design enables coarse-to-fine discretization of grasp parameters, encouraging semantic separation between global motion, articulated finger pose, and fine-grained refinement. The joint token $\langle j\rangle$ is incorporated as conditioning information to ensure articulation-aware manipulation modeling.
[…]
This structured decoding explicitly models hierarchical token dependencies, allowing articulation states to influence global motion, articulated pose, and fine-grained refinement in a unified manner. As a result, the discrete representation captures object articulation dynamics and coordinated hand manipulation within a semantically aligned latent space.
Figure 3: HAOI Manipulation Language Model
3.3 HAOI Manipulation Language Model
Benefiting from the previously introduced Discrete Articulated Manipulation Representation, we further develop a HAOI Manipulation Language Model. As illustrated in Fig. 3, our model is designed to align three heterogeneous modalities: human manipulation, object features, and language instructions.
After training the VQ-VAE, its parameters are frozen. We then utilize the VQ-VAE encoder $\mathcal{E}$ to obtain discrete tokens from HAOI sequences. The specific tokens provided to the MLLM vary depending on the downstream task. For the HAOI generation task, the objective is to predict subsequent grasps conditioned on the current object state; therefore, only the current object joint token $\langle j\rangle$ is required as input. In contrast, the HAOI prediction and interpolation tasks depend on the complete state of the current or intermediate interaction, and thus require the full token sequence $\langle g\rangle\langle l\rangle\langle r\rangle\langle j\rangle$ as input. To distinguish manipulation tokens from textual instructions, special markers <SG> and <EG> are appended to the token sequence to explicitly denote its beginning and end.
Regardless of the downstream task formulation, the generation objective of the MLLM is to produce a complete HAOI sequence, namely $\langle g\rangle\langle l\rangle\langle r\rangle\langle j\rangle$. To further differentiate generated manipulation tokens from natural language tokens, we introduce special tokens <HO> to mark the start and end of manipulation generation. The overall generated grasp sequence can thus be expressed as follows:
$\small\langle HO\rangle\langle g_{1}\rangle\langle l_{1}\rangle\langle r_{1}\rangle\langle j_{1}\rangle\cdots\langle g_{t}\rangle\langle l_{t}\rangle\langle r_{t}\rangle\langle j_{t}\rangle\langle HO\rangle,$
where $t$ is the total number of steps. For HAOI manipulation generation, $\langle HO\rangle$ tokens are removed, and each tuple $\langle g,l,r,j\rangle$ is sequentially mapped through the pre-trained multi-layer VQ-VAE codebook to obtain latent vectors. These vectors are then decoded by the VQ-VAE decoder $\mathcal{D}$ to produce the grasp configuration for each frame.
$P(x_{i}|x_{<i})=P(x_{i}|x_{1},x_{2},\ldots,x_{i-1}),i=1,2,...,t.$ (1)
3.4 Articulation-Aware Training Objective
We train the proposed framework in two stages: (1) learning discrete articulated manipulation representations via VQ-VAE, and (2) fine-tuning the HAOI manipulation language model. The overall objective integrates geometry-aware regularization, hierarchical reconstruction, vector-quantization commitment, and sequence modeling losses.
Geometry-Aware Regularization. To ensure physically plausible hand–object interactions and accurate joint-state modeling, we introduce two geometric constraints.
The penetration loss penalizes interpenetration between the predicted hand mesh $\hat{\mathcal{H}}$ and the articulated object mesh $\mathcal{O}$. Let $V$ denote the vertices of $\hat{\mathcal{H}}$, and $V_{\text{in}}\subseteq V$ be the subset penetrating the object surface. For each $v_{i}\in V_{\text{in}}$, we compute the squared distance to its closest surface point $v_{i}^{\mathcal{O}}$:
$\mathcal{L}_{\text{pen}}=\frac{1}{|V_{\text{in}}|}\sum_{v_{i}\in V_{\text{in}}}\left\|v_{i}-v_{i}^{\mathcal{O}}\right\|_{2}^{2}.$ (2)
For the articulated object branch, the joint reconstruction loss ensures accurate recovery of the joint state $J$:
$\mathcal{L}_{\text{joint}}=\left\|\mathcal{D}_{j}(\hat{z}_{j})-J_{\text{gt}}\right\|_{2}^{2},$ (3)
where $\hat{z}_{j}$ denotes the quantized latent representation of the joint branch. The geometry-aware regularization is thus:
$\mathcal{L}_{\text{geom}}=\lambda_{\text{pen}}\mathcal{L}_{\text{pen}}+\lambda_{\text{joint}}\mathcal{L}_{\text{joint}}.$ (4)
Hierarchical Reconstruction Loss. For the multi-stage VQ-VAE modeling hand manipulation, reconstruction is performed at three hierarchical levels corresponding to the global ($g$), local ($l$), and refinement ($r$) branches. Let $\mathcal{M}(R,P,T)$ denote the MANO mapping from grasp parameters to mesh vertices. The three levels are unified into a single reconstruction objective:
$\begin{aligned} \mathcal{L}_{\text{rec}}=&\;\lambda_{g}\left\|\mathcal{M}(R,0,T)-\mathcal{M}(\hat{R},0,\hat{T})\right\|_{2}^{2}\\ &+\lambda_{l}\left\|\mathcal{M}(R,P,T)-\mathcal{M}(\hat{R},\hat{P},\hat{T})\right\|_{2}^{2}\\ &+\lambda_{r}\left\|\mathcal{M}(R,P,T)-\mathcal{M}(\Delta\hat{R}\cdot\hat{R},\Delta\hat{P}+\hat{P},\Delta\hat{T}+\hat{T})\right\|_{2}^{2}\end{aligned}$
The first term supervises coarse global pose prediction, the second term refines articulated hand configuration, and the third term enforces residual correction for fine-grained alignment.
VQ-VAE Commitment Loss. For each branch $i\in\{g,l,r,j\}$, the encoder output $z_{i}$ is quantized to its nearest codebook embedding $\hat{z}_{i}$. The commitment loss stabilizes codebook learning and prevents encoder drift:
$\mathcal{L}_{\text{com}}=\sum_{i\in\{g,l,r,j\}}\left(\left\|\mathrm{sg}[z_{i}]-\hat{z}_{i}\right\|_{2}^{2}+\beta\left\|z_{i}-\mathrm{sg}[\hat{z}_{i}]\right\|_{2}^{2}\right),$ (6)
where $\mathrm{sg}[\cdot]$ denotes the stop-gradient operator.
Manipulation Language Modeling Loss. For the HAOI manipulation language model, sequence generation is formulated as next-token prediction. Given token sequence $\{x_{i}\}_{i=1}^{t}$, the negative log-likelihood loss is:
$\mathcal{L}_{\text{NLL}}=-\sum_{i=1}^{t}\log P(x_{i}\mid x_{<i}).$ (7)
To enforce temporal consistency between adjacent frames, we introduce a pose consistency loss:
$\mathcal{L}_{\text{temp}}=\begin{cases}\sum_{i=1}^{n}\arccos\left(\frac{\mathrm{Tr}\!\left(\Delta\mathcal{R}_{J_{i}}\Delta\mathcal{R}_{R_{i}}^{\top}\right)-1}{2}\right),&\text{rotational joint,}\\[6.0pt] \sum_{i=1}^{n}\|\Delta J_{i}-\Delta T_{i}\|_{2},&\text{translational joint.}\end{cases}$
Based on the decoding procedure described in Section 3.2, the output tokens of the manipulation language model are converted into the joint axis state $J$ for each frame. For rotational joints, $J$ encodes the object’s axis–angle representation. We convert $J$ into a rotation matrix $\mathcal{R}_{J}$ using the Rodrigues formula, and compute the inter-frame rotational change as $\Delta\mathcal{R}_{J_{i}}=\mathcal{R}_{J_{i}}\mathcal{R}_{J_{i-1}}^{\top}$. The hand’s rotational change $\Delta\mathcal{R}_{R_{i}}$ is obtained analogously. For translational joints, $J$ directly represents the object’s translation vector; thus, $\Delta J_{i}$ and $\Delta T_{i}$ denote the inter-frame displacement of the object and the hand, respectively. Here, $n$ denotes the number of inter-frame transitions, excluding the first frame which has no predecessor.
Overall Objective. The final training objective jointly minimizes:
$\mathcal{L}=\mathcal{L}_{\text{geom}}+\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{com}}+\mathcal{L}_{\text{NLL}}+\lambda_{\text{temp}}\mathcal{L}_{\text{temp}}.$ (9)
We collectively refer to the above formulation as an articulation-aware learning objective, as it explicitly enforces geometric validity, joint-state correctness, hierarchical grasp reconstruction, and temporal articulation consistency.
3.5 HAOI-Lang Dataset Construction
@@
To ensure consistency in data generation, we first assume that each object contains only a single type of joint, while allowing for multiple joints of that type. Furthermore, each joint type is associated with a predefined operational range of motion. The joint categories considered in our setting include the two most common types: revolute and prismatic joints. For any grasp point p{p} relative to the object’s center of mass $g_{\mathcal{O}}$, the grasp direction is ${D}={p}-{g}_{\mathcal{O}}$. Each frame provides a homogeneous transformation $\mathcal{T}\in\mathbb{R}^{4\times 4}$; we normalize coordinates by applying $\mathcal{T}^{-1}\mathbf{x}$, mapping all grasping data into a unified canonical space.
@@
To capture human intent, we render GraspXL grasping sequences in Open3D and prompt GPT-4 to produce multi-perspective textual descriptions covering contact points, wrist orientation, and intended action. This ensures concise, consistent annotations (Figure 4). Our dataset comprises seven object categories, 256 object instances, and over 50,000 manipulation sequences in total.